OpenStudy (anonymous):

Solve for x: (x+4)/3= 2 x = −2 x = 2 x = 2/3 x = -10/3

OpenStudy (goldphenoix):

\[\large \large \frac{ x+4 }{ 3 } =2\] What do you think you do first?

OpenStudy (anonymous):

divide

OpenStudy (goldphenoix):

Hmm. What do we divide?

OpenStudy (anonymous):

the x value

OpenStudy (goldphenoix):

So what do you get?

OpenStudy (anonymous):

-2

OpenStudy (goldphenoix):

I don't get what you did. o.o. So what did you do?

OpenStudy (anonymous):

@GoldPhenoix i think you might be leading him in the wrong direction

OpenStudy (goldphenoix):

I want to see his mistake first. :) So I can help him fix his mistake.

OpenStudy (goldphenoix):

This is how you find x, you want to multiply 3 and 2. Why? Well because it would be easier to find x, and it's possible. So 3*2 = 6. So your equation should look like: \[\large \large x + 4 = 6\] Make sense?

OpenStudy (anonymous):

so its 2

OpenStudy (goldphenoix):

Yes. X is equal to 2. How to make sure if it's 2? Well plug it in.\[\large \large \frac{ 2+4 }{ 3 } =2\] The numerator becomes 6. 6/3 = 2. Yes, x = 2. (Once again) I hope this help!

OpenStudy (goldphenoix):

When you have an equation with fraction, and you want to solve for x. You want to multiply the base by the denominator. Alright?